1 ton of CaCO3 is used in a wet lime slurry process to remove 90% of the sulfur from 4 metric tons of coal, which is 2% pyrite. The product of this scrubbing process is CaSO4. What percentage of CaCO3 is converted into CaSO4?
Answer #1

Mass of pyrite in the coal = 2 / 100 * 4000 kg = 80 kg

Pyrite is FeS2. Mass fraction of S in pyrite = mass of S / total mass = 2 * 32 / (2 * 32 + 56) = 0.533

Mass of S in the coalc = mass of pyrite * mass fraction = 80 * 0.533 = 42.64 kg

90% of S = 90 / 100 * 42.64 = 38.38kg

Now, the equation of reaction =

4FeS2 + 15O2 + 8CaCO3 = 2Fe2O3 + 8CaSO4 + 8CO2

From this, we see moles of S = moles of CaCO3 = mass of S / molar mass = 38380 / 32 = 1199.375 mol

Mass of CaCO3 reacted = moles * molar mass = 119.375 * 100 = 11937.5 g = 11.938 kg

% reacted = mass reacted / mass initially * 100 = 11.938 / 907.2 * 100= 1.32 %